MsBrains (ms-brains):

Can someone please help me fill out this practice sheet? (Example, and question in comments)

MsBrains (ms-brains):

Question.

MsBrains (ms-brains):

Example, It should be filled out like this

MsBrains (ms-brains):

Graph

MsBrains (ms-brains):

Graph Example

MsBrains (ms-brains):

@FortyTheRapper

MsBrains (ms-brains):

I know the graph should be like this,

OpenStudy (fortytherapper):

For the first one, it says x = -4 So wherever you see a x in the original equation, you replace it with a -4 For the first one: \[y = \left| (-4)+2 \right| -1 See how it is basically the same equation, we just plugged in the given x value instead of writing x itself?

OpenStudy (fortytherapper):

\[y = \left| (-4)+2 \right|-1\]

MsBrains (ms-brains):

Oh! So like in the second column I would write; (-3)+2 -1 ?

OpenStudy (fortytherapper):

Exactly! Just remember the absolute value bars and you're good to go

MsBrains (ms-brains):

Okay, gimme 1 sec to write this down, and after, do you mind checking if I filled it in correctly?

OpenStudy (fortytherapper):

Sure, I dont mind

MsBrains (ms-brains):

How do I fill in the y, and (x,y) columns?

OpenStudy (fortytherapper):

(x,y) -- those are the values that you have on your graph. You can just copy those And remember to turn those parenthesis into straight lines. Those make a big difference in calculations

MsBrains (ms-brains):

Alright, just to be sure, is this how it should look?

MsBrains (ms-brains):

@FortyTheRapper You there?

OpenStudy (fortytherapper):

Oh yeah, I'm here mb. I'm working on some Physics myself And yep, its good! Just remember to add a "-1" in the equation part for x = -1 \[y = \left| (-1) +2 \right| -1\]

MsBrains (ms-brains):

Awesome! Thanks so much. I'm pretty sure I'll be able to the graph myself. You're the best! <3

OpenStudy (fortytherapper):

You're welcome, anytime!

MsBrains (ms-brains):

wait, I actually have one quick question, Doesn't the 'Show your work' column look a bit empty, are we supposed to add anything else there? @FortyTheRapper

OpenStudy (fortytherapper):

Nah, that's all that is necessary You could put y = in front of each one of those expressions, so they turn into valid equations

MsBrains (ms-brains):

Awesome. alright. Thanks again!